A motorist thought she had just enough cash to buy 12 gallons of gas. But the price was 30 cents per gallon higher than she had expected, so she could afford only 10 gallons. How much did gas cost per gallon?

Answer by CharlesG2(834)

A motorist thought she had just enough cash to buy 12 gallons of gas. But the price was 30 cents per gallon higher than she had expected, so she could afford only 10 gallons. How much did gas cost per gallon?
x expected price * 12 = cash
(x+0.30) * 10 = cash
12x = (x+0.30) * 10
1.2x = x + 0.30
0.2x = 0.30
x = 1.50
x + 0.30 = 1.80 which is amount gas cost per gallon